OpenStudy (anonymous):

Two factors of –48 have a difference of 19. The factor with a greater absolute value is positive. What is the sum of the factors? –19 –13 13 16

OpenStudy (anonymous):

@Nnesha

OpenStudy (perl):

Assume x is the greater number. You have a system of equations x * y = -48 x- y = 19

OpenStudy (anonymous):

Ok @perl So What Would iDo ?

OpenStudy (perl):

you can solve that by substitution

OpenStudy (anonymous):

C @perl

OpenStudy (perl):

sorry, I can't give out answer. But I can guide you

OpenStudy (anonymous):

"The factor with a greater absolute value is positive." This tells us that their sum must be positive.

OpenStudy (anonymous):

ok ! & iWas Just Asking Was That One Correct

OpenStudy (perl):

can you show your work

OpenStudy (anonymous):

writing it on paper

OpenStudy (perl):

ok :)

OpenStudy (anonymous):

do i just subtract ?

OpenStudy (anonymous):

No, solve any of the equations and substitute.

OpenStudy (perl):

in the second equation you can solve for x or y

OpenStudy (perl):

then plug this into the first equation

OpenStudy (perl):

or you can solve for either variable in the first equation and plug into second equation

OpenStudy (perl):

The first suggestion was to use the second equation, and let's solve for y So we have two equations, i labeled them (1) and (2) (1) x * y = -48 (2) x- y = 19 in (2) , solve for x by adding y to both sides x = 19 + y now plug this into (1)

OpenStudy (anonymous):

thank you :)